1. Beginning inventory, purchases, and sales data for tennis rackets are as follows: April 3 Inventory...

1.

Beginning inventory, purchases, and sales data for tennis rackets are as follows:

April 3 Inventory 21 units @ $17
11 Purchase 12 units @ $17
14 Sale 16 units
21 Purchase 14 units @ $19
25 Sale 17 units

Complete the inventory cost card assuming the business maintains a perpetual inventory system and determine the cost of goods sold and ending inventory using LIFO.

Purchases Cost of
Goods Sold
Inventory
Date Qty. Unit Cost Total Cost Qty. Unit Cost Total Cost Qty. Unit Cost Total Cost
April 3 $ $
       11 $ $ $ $
$   
       14 $ $ $ $
$   
       21 $ $ $ $
$   
       25 $ $ $ $
$   
Total Cost of goods sold $ Ending inventory value

$

-------

2.

Journalize the following transactions using the direct write-off method of accounting for uncollectible receivables.

April 1 Sold merchandise on account to Jim Dobbs, $6,900. The cost of goods sold is $4,600.
June 10 Received payment for one-third of the receivable from Jim Dobbs and wrote off the remainder.
Oct. 11 Reinstated the account of Jim Dobbs and received cash in full payment.

If an amount box does not require an entry, leave it blank.

April 1
April 1
June 10
Oct. 11
Oct. 11

-------

3.

Equipment was acquired at the beginning of the year at a cost of $78,660. The equipment was depreciated using the straight-line method based upon an estimated useful life of 6 years and an estimated residual value of $7,920.

Required:

a. What was the depreciation expense for the first year?
b. Assuming the equipment was sold at the end of the second year for $59,486, determine the gain or loss on sale of the equipment.
c. Journalize the entry to record the sale. Refer to the Chart of Accounts for exact wording of account titles.
Answer #1

Complete the inventory cost card assuming the business maintains a perpetual inventory system and determine the cost of goods sold and ending inventory using LIFO.

Purchases Cost of
Goods Sold
Inventory
Date Qty. Unit Cost Total Cost Qty. Unit Cost Total Cost Qty. Unit Cost Total Cost
April 3 21 $17 $357
       11 12 $17 $204 21 $17 $357
12 $17 204  
       14 12 $17 $204 17 $17 $289
4 $17 68  
       21 14 $19 $266 17 $17 $289
14 $19 266
       25 14 $19 $266 14 $17 $238
3 $17 51  
Total Cost of goods sold $589 Ending inventory value

$238

2) Journal entry

If an amount box does not require an entry, leave it blank.

April 1 Account receivable 6900
Sales revenue 6900
April 1 Cost of goods sold 4600
Inventory 4600
June 10 Cash 2300
Allowance for doubtful account 4600
Account receivable 6900
Oct. 11 Account receivable 4600
Allowance for doubtful account 4600
Oct 11 Cash 4600
Account receivable 4600

3a) Depreciation expense = (78660-7920/6) = 11790

3b) Gain (loss) = 59486-(78660-11790-11790) = 4406

3c) Journal entry

Date account and explanation Debit Credit
Cash 59486
Accumulated depreciation 23580
Gain on sale of equipment 4406
Equipment 78660